# Khovanov homology can distinguish exotic Mazur manifolds

### v2

@@ Abstract.
A Mazur manifold is a compact, contractible $4$-manifold that has a handle decomposition with a single $1$-handle and a single $2$-handle. We show that Khovanov homology can distinguish certain exotic Mazur manifolds.
@@ 1. Introduction
Ren and Willis [RW24] gave the first analysis-free proof of the existence of exotic compact, orientable $4$-manifolds; their main tools are the Khovanov skein lasagna module and the Khovanov skein lasagna $s$-invariant. We first deduce from Hayden and Sundberg’s theorems [HS24] that Khovanov homology can distinguish exotic compact, orientable $4$-manifolds (Corollary 2.3). Then, we use similar ideas to show that Khovanov homology can distinguish exotic Mazur manifolds, i.e. compact, contractible $4$-manifolds that have handle decompositions with a single $1$-handle and a single $2$-handle. The first pair of exotic compact, contractible $4$-manifolds was obtained by Akbulut and Ruberman [AR16], and examples of exotic pairs of Mazur manifolds have previously been discovered as well [AY19, HMP21].
@@
Theorem 1.1 is proved in Section 4. These $4$-manifolds are the exteriors of the disks in $\overline{\mathbb{CP}^{2}}\backslash{\rm int}D^{4}$ given by blowing up Hayden and Sundberg’s exotic asymmetric disks in $D^{4}$ (Figure 2.1) at a point on the disks: see Figure 4.1 for alternative handle diagrams of these manifolds. We thank Kyle Hayden for pointing out that they are Mazur manifolds.
Ren and Willis [RW24, Section 6.11] defined the Khovanov cobordism map for oriented surfaces in $\mathbb{CP}^{2}\backslash{\rm int}D^{4}$, and showed, using skein lasagna modules, that it only depends on the isotopy class of the surface rel $\partial$. A key ingredient of our proof of Theorem 1.1 is Lemma 3.3, which says that this cobordism map is invariant under diffeomorphisms of $\mathbb{CP}^{2}\backslash{\rm int}D^{4}$ that fix the boundary pointwise. Although we were motivated by skein lasagna modules, we also present a proof in Subsection 3.1 that avoids the theory of skein lasagna modules, with the aim of making the argument more transparent.
@@ Acknowledgements
We thank Peter Ozsváth for his continuous support and helpful discussions. The main observation of this work was made while rereading the author’s Part III essay during a visit to Duke University. We thank Duke University for their hospitality and Jacob Rasmussen for advising the author’s Part III essay Khovanov Homology and Embedded Surfaces. We thank Nathan Dunfield for help with SnapPy, Kyle Hayden, Adam Levine, Lisa Piccirillo, Misha Schmalian, Alison Tatsuoka, and Qiuyu Ren for helpful discussions, and Kyle Hayden and Robert Lipshitz for helpful comments on earlier drafts.
@@ 2. Exotic compact, orientable $4$-manifolds from exotic disks
Recall that the cobordism maps on Khovanov homology are well-defined up to sign, and only depend on the isotopy class rel $\partial$ of the surface [Kho00, Jac04, MWW22].

### v3

@@ Abstract.
In a recent breakthrough, Ren and Willis gave the first analysis-free proof of the existence of exotic compact, orientable $4$-manifolds; their main tool is the Khovanov skein lasagna module defined by Morrison, Walker, and Wedrich. In this paper, we introduce a new, simple way of using Khovanov homology to distinguish certain exotic compact, orientable $4$-manifolds; our new method does not depend on the skein lasagna module. As an application, we give the first analysis-free proof of the existence of exotic Mazur manifolds, i.e. compact, contractible $4$-manifolds that have handle decompositions with a single $1$-handle and a single $2$-handle.
@@ 1. Introduction
In a recent breakthrough [RW24], Ren and Willis gave the first analysis-free proof of the existence of exotic compact, orientable $4$-manifolds; their main tool is the Khovanov skein lasagna module defined by Morrison, Walker, and Wedrich [MWW22]. In this paper, we first give a simple, skein lasagna-free proof (Corollary 2.4) of the existence of exotic compact, orientable $4$-manifolds. This proof only uses that the Khovanov map for link cobordisms $S\subset[0,1]\times\mathbb{R}^{3}$ only depends, up to sign, on the isotopy class of $S$ rel $\partial$ [Kho00, Jac04], and Hayden and Sundberg’s proof that this Khovanov map distinguishes certain exotic disks in $D^{4}$ [HS24].
We then extend our method and give the first analysis-free proof (Theorem 1.1) of the existence of exotic Mazur manifolds, i.e. compact, contractible $4$-manifolds that have handle decompositions with a single $1$-handle and a single $2$-handle. This proof is also skein lasagna-free and uses little machinery; the key technical input is Ren’s partial computation of the Khovanov homology of torus links $T(n,n)$ and maps between them [Ren24]. Note that the first pair of exotic compact, contractible $4$-manifolds was obtained by Akbulut and Ruberman [AR16], and examples of exotic pairs of Mazur manifolds have previously been discovered as well [AY19, HMP21].
@@
Theorem 1.1 is proved in Section 4. These $4$-manifolds are the exteriors of the disks in $\overline{\mathbb{CP}^{2}}\backslash{\rm int}D^{4}$ given by blowing up Hayden and Sundberg’s [HS24] exotic asymmetric disks in $D^{4}$ (Figure 2.1) at a point on the disks: see Figure 4.1 for alternative handle diagrams of these manifolds. We thank Kyle Hayden for pointing out that they are Mazur manifolds.
Ren and Willis [RW24, Section 6.11] defined the Khovanov cobordism map for oriented surfaces in $\mathbb{CP}^{2}\backslash{\rm int}D^{4}$, and showed, using skein lasagna modules, that it only depends on the isotopy class of the surface rel $\partial$. A key ingredient of our proof of Theorem 1.1 is Lemma 3.3, which says that this cobordism map is invariant under diffeomorphisms of $\mathbb{CP}^{2}\backslash{\rm int}D^{4}$ that fix the boundary pointwise. Although we were motivated by skein lasagna modules, we also present a proof in Subsection 3.1 that avoids the theory of skein lasagna modules and instead uses [Ren24] directly, with the aim of making the argument more transparent.
@@ Acknowledgements
We thank Peter Ozsváth for his continuous support and helpful discussions. The main observation of this work was made while rereading the author’s Part III essay during a visit to Duke University. We thank Duke University for their hospitality and Jacob Rasmussen for advising the author’s Part III essay Khovanov Homology and Embedded Surfaces. We thank Nathan Dunfield for help with SnapPy, Kyle Hayden, Adam Levine, Lisa Piccirillo, Misha Schmalian, Alison Tatsuoka, and Qiuyu Ren for helpful discussions, Kyle Hayden and Robert Lipshitz for helpful comments on earlier drafts, and Selman Akbulut for pointing out a reference.
@@ 2. Exotic compact, orientable $4$-manifolds from exotic disks
Recall that the link cobordism maps on Khovanov homology are well-defined up to sign, and only depend on the isotopy class rel $\partial$ of the surface [Kho00, Jac04, MWW22].
Remark 2.2.
For Theorem 2.1, we do not need that Khovanov homology satisfies the sweep-around move [MWW22]; we only need that the Khovanov map for link cobordisms $S\subset[0,1]\times\mathbb{R}^{3}$ only depends, up to sign, on the isotopy class of $S$ rel $\partial$ [Kho00, Jac04].
Indeed, let $\Sigma,\Sigma^{\prime}\subset D^{4}$ be two disks such that $\partial\Sigma=\partial\Sigma^{\prime}$. Let $x\in\partial D^{4}$ be away from $\partial\Sigma=\partial\Sigma^{\prime}$, and identify a small closed neighborhood $\overline{N}(x)\subset D^{4}$ of $x$ with $0,1]\times D^{3}$, where $\overline{N}(x)\cap\partial D^{4}=0\times D^{3}$ and $x=(0,0)\in 0\times D^{3}$. If there exists an isotopy rel $\partial$ between two disks $\Sigma,\Sigma^{\prime}\subset D^{4}$, then the isotopy does not intersect a sufficiently small neighborhood $0,\varepsilon]\times\varepsilon D^{3}\subset\overline{N}(x)$ of $x$. Identify $D^{4}\backslash([0,\varepsilon)\times\varepsilon D^{3})\cong[0,1]\times\mathbb{R}^{3}$; then $\Sigma,\Sigma^{\prime}$ are link cobordisms in $0,1]\times\mathbb{R}^{3}$ which are isotopic rel $\partial$ in $0,1]\times\mathbb{R}^{3}$.